I went through the table today looking for ghost beams. Most were already dumped. For those that weren't, I put down a dump or an iris.
I again looked at TTFSS OUT2 with the cavities unlocked (i.e., the open-loop error signals) and found that the low-frequency seismic/scatter wall appears only on south. So I hunted around south for a while. I found a series of ghost beams reflecting off the EOAM input and hitting dangerously close to the EOM output aperture. So I moved the EOAM forward a few inches, then adjusted its kinematic mount to offset these beams a bit. The EOAM should be realigned, and we should check to make sure the ghost beams are not entering the EOM again.
With the increased space between the EOM and EOAM, I installed a flipper mirror that takes the beam to the 1811. Then I minimized the RAM (from –54 dBm to –72 dBm with 85 mV dc).
FM dev: 10 kHz
Averages: 10, 50, 100, 500
I added a flipper mirror before the north EOAM, as well as a HWP before the resonant EOM (so that we can independently control the polarization entering the two EOMs). I optimized the RAM, but saw no improvement in the beat.